Docstoc

Mosaic_Engineering_Status_10_20_2008_thru_10_26_2008

Document Sample
Mosaic_Engineering_Status_10_20_2008_thru_10_26_2008 Powered By Docstoc
					From: Stein, Jack
Sent: Sunday, November 02, 2008 5:44 PM
To: Johnson, Bob; Smelser, Ron
Cc: Gay, Earl; Iyanar, Lokeshvar; Mehta, Sushant; Mori, Jayesh; Palaniswami, Siddharth; Tikekar, Rewa;
Venkataraman, Anuradha; Castaneda, Armando; Desgraves, Sean; DeWitt, Charles; Dyer, Rodney;
Edgecombe, Jason; Rhodes, Terry; Simineau, Wes; Whisler, Dave
Subject: Mosaic Computing Status for week ending Oct 20,2008 – Oct 26,2008

Mosaic Computing Status for week ending Oct 20,2008 – Oct 26,2008
Jack Stein:
 Key highlights
     A Mosaic Windows emergency update was done based on an out of cycle critical fix from Microsoft. All
         MWS systems, PCS systems, Active Directories systems, and other Windows systems used in the Mosaic
        infrastructure were updated:
     We have started to look at the impact of increased applications install on our 80GB based Mosaic
        Windows systems. We have 570 Mosaic Windows systems with 80GB of hard drive and we are
        approaching this maximum with each new software product we release.
     We continue to address the problem with use of the Centra software and OneNote in the Woodward 140
        Lab when Dr. Ivan Howitt uses the system to conduct a distance ed class with WCU. A concerted effort is
        in place to replicate the problem. Rodney is coordinating the efforts with 2 TA students (Earl and
        Nathan). Have been able to replicate the problem without FS in the picture. Gave Dr. Howitt an work
        around. Plan to get the OneNote problem escalated to Microsoft.
     Working with Jay Dominic get twice per year PC buy/replacement process (instead of once per year). We
        have about 50 Mosaic Windows primary workstations that need replaced (from 2005). We are working
        with Dell to get the newly announced Optiplex 760 in here for evaluation.
     A non-Mosaic Windows PC is set up in CARC 235 connected to the network and running logged in to
        Groove. This enables all members of Dr. Gary Wilson’s class to upload and share documents.
     Continued to work on the weather station on the CARC roof and the acquisition of data from it so it can be
        displayed on our CARC 101 digital display. This is an initiative from Dr. Saunders of Civil.
     The effort by facilities to determine the cause and the repair of the CARC window water leakage has been
        awaiting resolution (mainly funding).
     Working to identify items we would need if one time money was available.
     We plan to shut down the Centra EMeeting and discard all our licenses as FcTel is not interested in the
         transfer all of our Centra licenses to them.
     We continue dealing with very large number of application install for the faculty. Some are not really
        appropriate for College wide deployment.
     Our testing effort with NCSU’s VCL has been successful and the first effort with installation of the
        IMPLAN software for Dr. Ertunga Ozelkan and a few of his graduate students has been completed. So far
        this has worked out well.
     Terry is continuing with Primary and Secondary workstation deployment for faculty and staff.
     Continue to monitor and track the plethora of projects across Mosaic, PCS and with Support.
     We are working to evaluate VMWare and the ESX module as the kernel (replacing Linux) to determine
        how an RWS system performs with many more Mosaic XP guests on the same box we currently only can
        support 2 guests using VMWare Server. In parallel, we are received a loaner Dell server (r805 with ESX
        in firmware) using iSCSI disk array for the purpose of determining how many virtual Windows XP guest
        images can be supported by this configuration. Benchmark testing by our students has been completed.
        Analysis of the data, fine tuning the parameters and a re-test is in progress.
     Work is in progress to build a Red Hat Linux 64 bit release by this Dec 2008.
     Preparing for the audit visit – set for November 5,2008
     Dan Rowe has met with me and he is requesting a dash board view of all workstation, server rooms, storage
        rooms as part of an audit of how much is for research and how much is for academic use. I’ll be working on
        this.
     Attended an ITMT meeting were projects across campus was reviewed. We got thru about 50% and will be
      finishing this up in Nov 2008. I found out about some projects that would be beneficial to the COE and am
      following up with Johnna Watson to get more information.
     Attended a Media Storage meeting to review what we would need to from the central IT group for storage
      of media such as movies.
     Met with Howie Fang on his views about large Disk space offerings when a project ends; Research
      Computing to COE researchers from Mosaic.


  Customer Support:
   Had a meeting with the support team to review issues, projects.
   Armando is working to get Duke 242 set up for 7*24 availability. We need to install web cams in this
      room.
   Armando Castaneda and Sean Desgraves continues working with the team on support issues.
   Armando is continuing to work on a project to monitor departmental and research Mosaic labs on a
      different SLA using Nagios, Lab Patrol and a web cam. The SEEM lab in Colvard 2046 is being
      monitored on a pilot basis. A SLA was drafted for SEEM and we are monitoring this departmental lab.
   Armando continues developing and shaping the Mosaic Help Desk and Mosaic TA staff. TA are writing
      How To pages on products.
   Mosaic Incidents Tracking for the week (these are ones logged in our incident tracking system, we handle
      many other incidents that are not logged and are resolved quickly):
                Total Incidents = 1718 Open/Waiting=63 Closed=1655

Infrastructure and Services:
        The disk farm service was announced on August 25. The disk farm will provide COE faculty
           researchers a place to store large data sets (200GB – 2 TB), with back up support, on our AFS Mosaic
           file system. Dave Whisler build the data base for this service. Howie Fang is using the service. No one
           else has come forth requesting this service. I plan to announce the availability of this service to get
           some more interest
        Terry and I are meeting with the High Capacity Work Group headed by Chuck Price. At this point the
           COE is the only place where large disk space with back up is available as a service. Other groups are
           trying to get ITS to provide the same service. Chuck Price is preparing a report on this initiative.
        The 5 R300 Dell Servers are in. 3 test machines for Jason / TS. 1 Machine for VMWare
           VirtualCenter. 1 Machine for TS test LAMP / Wiki. We are planning on displacing the 2 745’s in the
           CARC server room once these are installed. The 745’s can be used to refresh the Mosaic TS machines
           in CARC

Personal Computer Support:
      o I am working with Apple to make the MacBook $999 offering as the entry level laptop. We are also
           working with Apple on a holiday laptop promotion.
      o I am working with Lenovo to address several issues with the laptops we are recommending. Mike will
           be working with me on a holiday promotion.
      o I am meeting with Dell to come up with a holiday promotion.
      o Continued effort to work on the buildup of the Personal Computing Support group.
      o We have been working with Dave Whisler to make enhancements to the Work Order data base for the
           SOC initiative. Also working with shop inventory system to handle loaner laptop and projectors.
      o The phone number for the Personal Computer Support is x73355. Office location is in Smith 202a.
      o Armando Castaneda, Ian Pugsley, Sean Desgraves and I are meeting regularly and reviewing the SOC
           project and initiatives.
      o Dell Statement of Work ( SOW) arrived Jesh from legal gave me his feedback.
               o PCS Work Orders Summary Report
                                             PCS Work Orders Summary Report
                         WO's opened between                     07/01/08    and          10/26/08
                         Opened by Month
                               July 2008                                                  16
                                                           % Total Opened             13.33%
                               August 2008                                                23
                                                           % Total Opened             19.17%
                               September 2008                                              52
                                                           % Total Opened             43.33%
                               October 2008                                               29
                                                           % Total Opened             24.17%
                                                            Total Opened                  120
                          Last Worked by Month
                               July 2008                                                   14
                                                       % Total Last Worked           11.67%
                               August 2008                                                 21
                                                       % Total Last Worked           17.50%
                               September 2008                                              56
                                                       % Total Last Worked           46.67%
                               October 2008                                                29
                                                       % Total Last Worked           24.17%
                                                         Total Last Worked                120
                          Closed by Month
                               July 2008                                                  14
                                                           % Total Closed            12.07%
                               August 2008                                               21
                                                           % Total Closed            18.10%
                               September 2008                                            56
                                                           % Total Closed            48.28%
                               October 2008                                               25
                                                           % Total Closed            21.55%
                                                            Total Closed                 116


 Digital Signage:
             I am currently working with Mike Hermann and am managing the insertion of the contents in the
             Nexus driven Players.
          The Digital Signage with Nexus is set up in Hawthorne. It appears that I do have access the
             Hawthorne Mac Mini via vnc. Working with Cathy Blat on contents and access to the building for me.
          Working with Mike Herman to compile contents for ECE to use on the Digital Signage.
          An initiative is underway to have all UNCC based contents available for the digital signs.
          I am working with the new NCS software as the new product we will be using. I created a test plan for
             the product.
          I am training Rewa and Lokesh when time permits on the Nexus product.


Classroom Lecture Capture and Distance Education:
         I am working to support Tony Brizendine and Steve Clark on the Distance Education project for ET in
            Smith 202. We have made good progress with the PanOpto Coursecast 2.0 software. We have started
            to do testing with one of the video cameras we ordered; we expect the second camera to be available
            week of Oct 20 and we will test with 3 cameras (Canon zr900_. We are also testing integration with
            Blackboard. I provide Tony with weekly updates. I held a presentation to ET faculty on the status of
            the project. Steve and I are scheduling a status meeting for Oct 29 for a more widespread group of
             people to determine what is the campus position on making the Panopto lecture capture go from testing
             into production.
            Working on SEEM Lecture Capture initiative. They are interested in Panopto for later when it is
             ready.
            Harish Cherukuri will migrate to PanOpto when it is ready.

 WEB Projects:
    . Sid an Anu are working on the new MEES web site pages to make sure it meets all the requirements of
      ME and the graphic designer.
    The Mosaic Web site is being re-assessed and we will rework the menu and contents to make it more
      intuitive and easier to find things.
    We plan to start testing another wiki open source product as the xwiki is not working out for us in our
      extensive testing efforts on the Nexus project.
    We will be working with Larry Sharp to set up ASME as a virtual web site.
    We plan to meet with the Motorsports web developer that created the web site for Dr. Jim Cuttino
       (motorsportu.com)
    Project to re-do the web form for appeal by suspended students (for Karina Franco) has been released for
      Karina to test and assess.
    Work is in progress to update LiveHelp web application (we are about 2 years behind). It was released but
      later on we had to pull it back due to issues with the MySQL server.
    Work is in progress to update the Web calendar application. It was tested and is ready for release.
    Work is in progress to evaluate use of YouTube for placement of our videos
    Work is in place to have our COE web news be available as RSS feed.
    Work has started to create a web based application to manage our PCS inventory (i.e. loaner laptops,
      projectors, etc), our consumable (toner, paper, etc), guest accounts, and vacation planner.
    ITS (Carter Heath) has blocked the release of IP addresses for our 2 virtual host systems. He is requesting
      that we update some of our software components on the COE Web but we will plan to do that in Dec
      2008. I am working with Jason on getting us new SSL certificates for the 2 secured COE web sites:
      https://www.coe.uncc.edu https://secure.coe.uncc.edu when we get these and Jason has these in place on
      our COE web server I will announce to all of COE. Right now when accessing these sites we are
      recommending to press continue when getting a warning message.
      Hit Counts on our existing Virtual Tour for COE: Since May 4,2007: 3692

Miscl:
        We continue to use Nagios open source web based software to monitor all our services in real time. The TA
         are also using this tool to monitor any issues with our services.

New Initiatives:
              We are researching the Wake-on-Lan feature that will let us turn on Dell Windows PC systems
                 that are turned off. Smith 260 is configured this way. With this feature we can now go “green” by
                 shutting down our Mosaic workstations (remotely) in labs that are closed and then turning them on
                 when we open a lab.
              Wes worked with Steve Clark’s group to deploy an elegant method of logging off a user who
                 forgot to do so in a podium. Deployed in Woodward 140 and Alan Hege has no issues with it.
                 Deployed in Smith 21,218,219 podiums and Help Desk to test. We plan to release this to all
                 Mosaic based podiums.
              The virtual Windows XP system has been released as Beta to Brian Anderson and Harish
                 Cherukuri as well as Ron Smelser. A User Guide showing the installing and managing of this
                 offering has been produced. So far Harish and Ron have been very enthusiastic with this product.
                 Harish has found some performance issues with AFS access and we have diagnosed the issue and
                 provided a fix that appears to have fixed the issue. We are working to get a release update so we
                 can seed Tony Brizendine and Steve Patterson. Work is in progress on VMWare and PRS
              The Kodak digital picture frame (10”) with wireless support is installed in Smith 202a (PCS
                 Help Desk) We received and are testing the picture frame with direct Ethernet support (as well as
                   wireless) from Dlink. We plan to test this to see if we can use it for Mosaic labs/classrooms to
                   show room schedule and other content.
                  PDA strategy for the COE. We seeded iPod Touch to Bob Johnson, Ron Smelser, Tony
                   Brizendine, Dr. Chung-Suk Cho, Martin Kane, Bill Heybruck, Yogi Kakad, Helene Hilger, Bev
                   Guessford, Jack Stein, Patty Tolley. At this point the project is complete.
                  Looking at a different methodology for event calendar on COE Web.
                  Looking at a different methodology of how room reservation are managed. Looking at wireless
                   digital photo displays mounted on wall or using of Exchange shared calendar for the handling the
                   rooms (this is how HSS and C of Ed does it). Also looking at R25 as another possibility.
                  We will be testing the Linux OS as the file server OS – replacing over a 2-3 year period the Sun
                   OS based file servers.
                  Wes has been migrating the license services off ls-sm1 to ls-me1

Applications:
    We plan to purchase Adobe CS-4 in Nov 2008 and install it on Mosaic.
    We are working on the Abaqus 6.8 on Linux for Harish
    Plan to meet w2ith Jim Conrad about Solidworks
    We plan to remove Eudora from Mosaic.
    Working on the Prolog software for Dr. Chung-Suk Cho of ET. We are running into difficulties with this
         package.
    ME has decided to fund the renewal of Adam software and we are waiting for the paper work.
    Working with Dr. John Gresser on PCB design software that is not in the Cadence suite but may be with
         Multisim
    Working on the software install for Dr. Mesbah Uddin: STAR CCM+
    Working to get software for Dr. Mesbah Uddin: CFD-ACE
    We have re-installed the Synopsys software on Unix and Linux. ECE will find an alternative product
         solution for their research and teaching needs. Work is in place with Ron Smelser and ECE to get sign off
         to remove splash page.
    No activity on the Aveva software. Rodney has been looking for cycles to get up to speed on this
         configuration. Areva has not been responding to any request for assistance. Rodney and I met with Ian
         Piggott of Aveva to see if he can help us kick start the PDMS installation. The ball is really in the
         departments to decide the classes and syllabus that incorporates this product.


Software purchase, renewal status:

           Expiration      Software                    Costs                Notes

                  6/1/08   GRL Weap                              Contacted Brian Anderson if it needs renewed; he needs
                                                                 to let us know
                 7/31/08   Respondus                             Chad needs to get updated version from FcTel
                 8/31/08   Algor                                 Need to check with Dr. Chen – he renews this.
                  9/6/08   ADAMS                        $1500    Working with Dr. Cuttino
                 12/1/08   Labview                      $2500    Being handled by Chuck Price

Planned for Nov 2008 Update:
        Microsoft Windows patches and Office patches

In the queue
          Windows Update Agent (prerequisite)
         Firefox 3.0
         PRS Clicker Software
         ABAQUS 6.8
          SmartBoard
          TNT Identity Client Removal
          JInitiator Removal (multiple versions)
          Origin 8.4
          ArcGIS 8.3 Removal (need to begin communicating this packages removal to faculty)
           AFS Client / Kerberos Client (assuming testing goes well)
           Oracle 11g – In Queue
           Visual ModFLOW Pro 4.3 – Low Priority
          P-Spice / OrCAD / Layout suite replacement
          Prolog
          Maple 12 on all 3 platforms
          PanOpto Coursecast Recorder and Editor software for lecture capture


On Hold
          Acrobat Pro
           Timberline Office – Current version is functioning postponing until possibly next summer
           Hyperworks 9 – Pushing out until either Winter break or next summer.
                            License issue driving update has been resolved.

We are considering using the ITS Citrix service for packages like SAS/JMP as well as ArcGIS.




Chad DeWitt:
   -   Worked with Wes to finalize FireFox 3.0.3 installation
   -   Fixed rws12 with login timeout issue
   -   Rebuilt VMWare Fusion machine, mws5, for Dr. Bowen
   -   Upgraded the memory in VMS-SM3 from 4 GB to 8 GB with Wes
   -   Tested emergency patch with Wes and Rodney
   -   Downloaded and packaged Panopto Recorder
   -   Downloaded and packaged Panopto Editor
   -   Performed patch cleanup
   -   Worked with Wes on getting the camera for the lecture capture project working
   -   Worked with Terry and Wes on mws7 issue

PTO
   -      December 19, 22, 23 (Please note change to include the 19 th)
   -      January 2


Wes Simineau:
---------------------
Following up on ProE issue in Duke 242 – Issue is resolved
Generating usage data for Jack on Maple
Provided status on ProLog software
            Requested clarification from vendor on software issues.
Worked with Chad on finalizing Firefox 3 package
Worked on building package for Pi Toolbox
            Built initial setup script for package
            requested and populated production volume
            Added links to test menu and sponsor to test group
            installed and tested package as user.
            Installed Package for testing by Dr. Tkacik
Worked on misc App cleanup with team.
Requested assistance from Jack on verifying suitability of new PRS software prior to adding to workload.
          Utilizing PCS staff / VMWare Player image
rebuilding / Updating SM3 test machines and machines in C228.
Misc updates to AppDB to reflect last update as well as new software requests.
Worked with Chad on debugging Canon camera for Lecture Capture project.
Followed up with Chris Maynard on keyboard / mouse issue.
Attended meeting with team / Dr. Fang on research computing needs.
Assisted Jason with debug of networking on new r200 Chassis in CARC server room
Updated all license servers for File sharing vulnerability
Scheduled / Monitored Thursday’s update for file sharing vulnerability of Mosaic XP machines
Worked with Rodney / Chad to test Mosaic XP emergency update.
Worked with Carter to resolve off-campus connectivity issue on Hawthorne display
Worked with Carter to setup SSL VPN for testing connectivity to F/S workstations from off-campus
Tested SSL VPN during update
Adjusted access list for UTEXASED4 software
Worked with Terry/Chad on issue with MWS7
Provided updates to Jack on Mosaic Infrastructure
Provided feedback to Jack on Budget
Provided information on MiKTeX problem for Jason
Managing / revising software application workload
Worked with Chad to swap memory in VMS-SM3 (upgraded to 8GB from 4)
Worked on Remote desktop issue for end user.


PTO
   -     December 22, 23
   -     January 2


Jason Edgecombe:
----------------------
Done:
* Sean & I spoke with Dr. Adams about what PDA to buy with his personal money. Recommended iPod touch.
* helped Dr. Nasipuri w/MikTeX
* Helped Nan Bousaba w/ WebPE
* helped Mike Martin to set working dir w/LS-Dyna
* fixed hspice/synopsys license error for Dr. Mukherjee
* fixed WebHW error at login screen by tweaking the code - no issues reported
* reported to Stantec that their company was added to the list of employers for COE grads.
* helped Dr. Adams with ssh timing-out overnight.
* created PCS TA's pts group and gave it access to vmplayer images. Gave Sean instructions on how to edit the
group/
* resolved an issue with new TA accounts where they their environment was not set up properly -- missing ta_bin
symlink in homedir.
* changed AFS cell server to restart at 4:30AM instead of 4AM to reduce noisy emails from failed cron jobs.
* closed issue where Andrew Willis had to reconfigure Matlab on his standalone machines to point to the new
license server.

Waiting:
* Waiting on Dr. Mukherjee for Wind River info.
* Dean Tolley will use a WebWork class for Fall 2008 that is hosted by the Math Dept. We will revisit in Dec 2008
to see if we need to have our own server.
* point cadence & pspice on solaris/linux to new license server

In Progress:
* working on SSL name error on www.coe SSL web sites with two-level names - blocked by ITS security until we
upgrade apache+php
* as part of 64bit Linux testing, test each Mosaic application on 64 bit and make an upgrade strategy.
* Started working Redhat 5.2 64-bit build -- got a box to build with 64bit RHEL5U2.
* helping Chris McDaniels with Webwork courses
* get R200 build with Linux for Wes -- will try 64bit. - requires drivers in RHEL5U1 or higher.
* AFS testing with outlook and OST files - installed latest client build
* investigate dropped packets from afs fileserver under load -- happens w/Outlook OST
* Earl: Installed Comsol (32/64) on 64bit Linux. -- waiting on testing
* nss_db causes gnome login to take too long on 64bit linux. Incident filed with RedHat.
* Earl installed on Solaris: ProENGINEER WildFire 4 -- waiting on testing.
* Earl installed Abaqus on 64bit Linux -- needs testing
* Earl installed Star CCM+ on 32bit Linux for Dr. Uddin - waiting on testing.
* working w/Dr. Soliman to help him access the ~saesb web page for editing.
* Earl installed Cadence5 on 64bit Linux -- needs testing, waiting on Weldon & Ravindran to settle on config files.
* helped Yi Yang w/Cadence 5 issue -- increased quota from 2GB to 8GB to see if his jobs still fail.
* setting up a test lxs server for 64bit testing.

In queue:
* Nagios work - needs extra custom plugins
* Linux as a file server project.
* upgrade php and apache on web server.
* reorganize/remodel shared ME folders for Robin Coger -- after Fall starts
* install on Linux/Solaris: ADAMS
* upgrade thunderbird on solaris
* upgrade thunderbird on linux
* troubleshoot why 32bit Maple 10 doesn't work on 64bit Linux.
* fix DNS server issues -- upgrade dns software
* add AFS rxdebug "calls waiting" metric to nagios
* fix apache logs archiving script to handle new ssl config
* shutdown samwrap license daemons on ls-sm1
* upgrade opnet on solaris/linux
* upgrade comsol on solaris
* install a better PDF reader on solaris - requested by Wes
* tweak alert timeouts in nagios for MosaicAnywhere machines to 2 hours. Add cron job to alert every 6 while in
maintenance.
* install Cadence6 on Linux
* install CCM+ on 64bit Linux for Dr. Uddin
* install CFD-ACE on Linux for Dr. Uddin
* setup www.asme.uncc.edu vhost on coe-web by end of Fall 2008 for Larry Sharpe.
* make a script that alerts Jack every hour if X+ number of rws machines are down.
* evaluate if Joshi's Precision 670's can run 64bit Linux
* Earl: installing AFS filedrawers software.

Postponed until Summer 2009 or later:
* RT Work - a test install is running. Working on customization.
* put mysql server on different hardware from oracle.



Dave Whisler:
-------------------
Mosaic: Attended weekly staff meeting

LiveHelp:
       Students release update to LiveHelp over the weekend which unitentionally caused MySQL to consume
        majority of CPU on database server.
       Spent much time on Tuesday investigating the problem of MySQL Daemon was consuming 95% of the
        database CPU.
       Analyzed the MySQL slow query log . Several LiveHelp queries were causing the CPU consumption.
        Sent them to the Students TA’s to fix.
       I shutdown the LiveHelp app until they were able to back-out their LiveHelp changes and then restarted
        LiveHelp App.
       Student TA’s are working on a way to correct the updates and they are to bring any future app changes to
        me before they go onto production.
       I have also provided the production MySQL config file for them to apply to their test server to more closely
        mimic the production system.

AdminDB:
    Worked on a long list of enhancements for Toby, below are just some of them.
         o Modifications to the Cost Sharing Obligations reports.
         o Combined a couple of separate reports into one report using run-time parameters to make it more
            flexible.
         o Will add a new table and several new fields to the database.
         o Will modify the front-end forms to handle the new table and new fields
         o Removed the SSN field from the AdminDB to comply with University policy (even though we
            had removed SSN data years ago).
         o Modifications to forms, tables, view, oracle packages and procedures to remove code involving
            ssn.
         o Removed the sec_level field from the admindb_sec table.
         o Removed code in the security procedure that utilized the sec_level field.
         o Consolidated 4 Student Appointments/Fellowships Reports into one report.
         o Removed other reports and cleaned up Reports form.
         o Fixed the dump to excel form for Salary appts by removing obsolete reference to SSN field.
         o Widened a field on the Fund Setup field
         o Removed an obsolete button on Reports Form and removed referencing code.
         o Majority of reports now have run-time options.
         o Looking into creation of a couple of new tables and fields to existing tables for Toby.

UserDB:
     Handled the normal Special Case, Guest account, Disk Quota increases and the other daily Mosaic user
        issues via PHP Support.

AppDB:
    Checked with Wes, Anu/Sid to make sure my recent changes for them are working.

Oracle:
     Assessing new Oracle CPU (Critical Patch Update) for October.

MRS DB:
    Looking into pulling info from Banner on all COE courses (name, number, section, instructor, schedule,
      building, room)
    Did research on capabilities of R25. Had good conversations with Donna Mclain and Lucian Wilhelm on
      its abilities for our use
    Had Lucian remove access to R25 from former COE staff.
    Met with Jack on this data need to help me design a proper solution.

PCS: Need to write documentation of MySQL back-end

MRS: Need to write documentation of MySQL back-end
SupportDB: Need to write documentation of Oracle back-end and Access front-end

Terry Rhodes:
•        Attended meeting with Howie Fang, Jack Wes and Rodney regarding Howie's
         needs and requests
•        Installed Mosaic PC in Woodward 467 Lab for Gloria Elliot
•        Installed new Mosaic PC DCH 221 for Gary Hodgins
•        Installed local printer in DCH 221 for Gary Hodgins
•        Installed local printer in DCH 238 (Gary's old office)
•        mws7 rebuilt (by Wes) because users could not login after update
•        Installed secondary Mosaic PC in CARC 254 for Environmental Office
•        Recovered T3400 evaluation unit from DCH 213
•        Reinstalled original Optiplex 755 in DCH 213
•        Processing RMA for bad memory from CDWG
•        Installed 4 R200 Dell servers in CARC 150B server rack
•        Registered 4 R200 servers for network access
•        Connected 4 R200 server to remote KVM and power and made necessary
         configurations changes to devices
•        Replaced defective machine in Smith 330 ET Shop
•        Working on surplusing more outdated equipment
•        Building volumes and mounts in AFS as requested
• Working to perform normal job duties
• Performed weekly AFS backups to tape
• Working to reorganize storage rooms

Vacation: Nov 24 - 26
           Dec 22 & 23
           Jan 2, 2009


Rodney Dyer:
-----------------
week ending 10/26/08:

* Meeting about disk farm.

* Post cleanup work after October Mosaic XP Maint update.

* Office work, email cleanup.

* Final migration of Mosaic AD services to new ADCSM3 server.

* Added Nathan Hatley (nhatley7) to needed groups for access to
  the CIFS directories (J:\coe\database\mosaic\supportdb,
  J:\coe\database\administrative\shop_inv).

* Cleanup of security access for Dave Whisler on J: drive
  security ACLs.

* Patching of the following servers after critical Microsoft
  SMB security hole problem was found...

 ADCSM1 (old account maintenance, deprecated)
 ADCSM3 (new account maintenance)
 ADCECE1 (AD fail over)
 ADCME1 (J: drive share, databases, outlook PST archives)

 MBS1 (Mosaic Build Server 1)

 MPS1 (Mosaic Probe Server 1, backup)
 MPS2 (Mosaic Probe Server 2, main)

* The following AD servers were backed up...

 ADCECE1 (AD fail over)
 MBS1 (Mosaic Build Server 1)
  + Update of OpenAFS client from v1.5.1 to v1.5.54.

* Emergency Mosaic XP Maintenance Update release of critical
  Microsoft patch WindowsXP-KB958644-x86-ENU for SMB security
  hole.

* Woodward 140 problems...

 Got Nathan and Earl going on debugging the OneNote issue.
 Helped Earl with problems related to his OneNote testing
 machine.

 Status of debugging:

 Nathan and Earl have been assigned testers under Rodneys
 direction.

 Nathan and Earl have built a non-Mosaic Windows XP machines
 with Office 2007 SP1.

 Nathan was able to replicate the OneNote crash without and
 apparently with mouse recording software.

 Earl is just getting up to speed with a properly configured
 XP machine for testing (will begin testing proper on Monday
 or Tuesday 10/27, 10/28).

 We are in search of a free, or possibly low priced purchase'able
 mouse recording and playback software for unit debugging. This
 should allow us to carry the step by step OneNote crash sequence
 from one machine to another. It should also allow us to change
 the configuration of the machine without requiring us to change
 the fault reproduction sequence.

 We are working to generate accurate data on the problem so that
 we can call Microsoft tech support with some level of confidence
 that we have done all that we can to prove the problem is with
 OneNote.

 We are deliberately working with different hardware and RAM
 configurations to establish that the problem is unrelated to
 hardware. Nathan has a fast Optiplex 745 machine with 2G RAM,
 and Earl has a slower GX270 with 1G RAM.

 For OneNote inking input, Earl is using a Wacom mouse input
    digitizing PAD, while Nathan is using a Symposium combination


Armando Castaneda:

     Informed Greg Kish again about labs not being clean
     Recovered missing flash drive for a student
     Closed Mosaic tickets
     Open Woodward labs
     Followed through for resolution to Jeff Raquet’s printer issue. ITS changed the printer server
      that the PC was using.
     Staff meeting
     Responded to Microsoft that a patch would be acceptable for the MathType issue
     Sent Larry Lane an email to schedule repairs on COE 3
     Assigned Earl Gay a training account (trng59)
     Update Lab reservation sheet
     Sent Rob Herrington an email concerning access to Lenel
     Sent email to Donna Overcash concerning the oversight in Eric Rhodes compensation
     Ran Office Diagnostics on Jim Conrad’s machine. It found and fixed 1 problem.
     Suggested the addition of user location to the ticketing system
     Took Raven Hayes (TA) out to lunch
     Created “How to” document for installing Windows patches
     Confirmed that Maple ver. 10 has a problem that ver. 12 does not. Informed student of options
     Received quote for Thin Client Kiosk to be installed in the Smith 2nd fl. common area
     Sent Dan Rowe an emails asking for information concerning our current lab alarm system

Sean Desgraves:
Weekly report for the week of: 10/20/08
•          Went to Staff meeting on Tuesday
•          Reviewed Camera Playback of 24 hour labs for non-business hours
•          Took various shifts in the PCS helpdesk to monitor and assist with customer issues and requests
•          Closed various Mosaic/PCS Tickets/WOs
•          Followed up with Kimberly Warren as to how the loaner is working out for Jacob Garbini
•          Setup a VMWare Windows XP image for testing – received/installed PRS software from Lee Perkins
•          Delivered iPod Touch to Helene Hilger
•          Delivered iPod Touch to Bev Guessford – working on multiple calendar problem with Apple
•          Checked out to Amy Kitts 4 laptops and a printer on Wednesday to be returned 10/27
•          Checked out to Yolanda McIlwaine 3 laptops on Friday to be returned 10/27
•          Picked up Robin Coger’s laptop (delivered to her by ITS) to install campus software – Closed WO
•          Delivered 2 Tablet PCs from ITS for Sharon Green/Churlzu Lim – contacted them with a list of software
to be installed
•          Contacted Patty Tolley again to schedule delivery of her iPod Touch – she’s “very busy”
•          Informed Jeff Fish in ITS (Desktop Support) of incoming MacBook for Wes
•          Assisted Kimberly Warren with a profile problem – Closed Ticket
•          Informed and followed up w/ Charis Mayhorn in ITS (Network Services) regarding various customer
systems registrations
•          Had PCS pick up 3 Dell Precisions from Dr. Joshi (2 setup for Mosaic XP and Mosaic Linux and 1 setup
as a standalone with Windows XP/Red Hat Linux dual-boot system)
•          Met with Jason and Ryan Adams for information on PDAs and the Apple iPod Touch
•          Provided initial report about Joseph Owen and Lenovo purchasing issues – Ticket No. 1822



Siddharth Palaniswami:
Last Week :
Worked on converting the graphic layout to HTML for MEES
Worked on making the HTML a template in Joomla
Showed Jayesh how to create panaromas from the images that Mike gives.
Documented the ticket system code and handed over to Sushant and Jayesh
Added How To to PCS page
Looked into LiveHelp problem
Attended COE Web Meeting
Updated PCS website with Apple flyer.


Anuradha Venkataraman:
I was working on the following last week:

Worked on converting the graphic layout to HTML for MEES
Worked on making the HTML a template in Joomla
Attended COE Web Meeting
Fixed attachment problem with editing tickets
Looked into LiveHelp problem
Added VCL and Citrix to S/W List and changed the Help File pop up and moved to production.
Showed Jayesh how to add Virtual Tours in Flash and created one as an example for Smith.
Made the Virtual Tours scrollable in beta.



Weekly Status report Sushant Mehta:
-----------------------------------------------------------------------------
Activities for week ending Oct 20,2008-Oct 27,2008

Summary:
1)Changes to suspension site.
2)Worked on the tab creation for Mosaic Account Activation site.
3) Worked on the Live Help database issue.
4) Tested and completed the Mosaic Account Activation site with ticket creation.
5) Worked on the ticket creation and emailing for additional quota request.
6) Understanding the ticket creation system.



Weekly report for of Jayesh Mori:

Weekly Status Report:

Monday:
> Understood the current incomplete labMap system and studied its code.

Tuesday:
> absent

Wednesday:
> Found the problem in livehelp and updated it.
> Learnt creating virtual tour from Anu and create one
> Groove new work space creation and invitation sending and confirming

Thursday:
> Running tests on live help to monitor impacts on database server
> Searched methods for implementing Drag and Drop

Friday:
> LabMap design and started coding.

Earl Gay:

Done:
- Finished copying applications from test volumes to production volumes.
- Deleted test mount points and cleared out unused volumes.
- Rebuilt LWS6.

Progress:
- Began working on installing AFS File Drawers; worked on setup and debugging issues.
- Discussed Nexus issue.
- Went to the CARC server room to help in setting up new servers.
- Installed Maple 12 on Solaris. (Waiting on testing)
- Installed Maple 12 for 32 and 64 bit Linux. (Waiting on testing)
- Menus and depot have been ran/created for Maple 12 on 64 bit Linux.
- Installed ADS 2008 and ICCAP 2008. (Waiting on testing)
- Continued researching dbus issue.
- Researched into HSPICE issues.
- Met with Mr. Dyer discussing debugging a OneNote issue.
- Set up XP SP3 and Wacom tablet and installed Office 2007 SP1 on LWS11.
- Set up automation with OneNote testing; tested with no error.
- Added version 2.0 of .NET to LWS11 and upgraded Graphics Driver.
- Continually met with Mr. Dyer working on OneNote issue.

Lokeshvar Iyanar:
Panopto lecture Capture:
 *We had recorded all the videos for Phase1 and Phase 2 with the new Canon camera. However, it is taking forever
to upload the DV lectures since they are of the size close to 3GB. The reason may be attributed to the network
traffic.

* I had uploaded all the lecture contents and updated version of the Panopto logbook herewith. Inferences had been
made and logged on all uploads.

* I had tried recording using two cameras by plugging them both to the Mac mini through the Fire wire hub. It did
not work out.

* I again tried recording by plugging them both to the fire wire ports fitted into the DELL desktop, but still the issue
did not get resolved.

Nexus Digital Sigange:

*I have understood on how to make a slide playlist and how to create a master playlist. I have created a new
presentation on the testing network and had created a master playlist. I had understood the whole process of how to
run a proper streamlined presentation in the LCD display. I understood how to make things go on and on in a loop. I
also learnt how to allot specific start and end times for specific slides.

* NEXUS - I had created the SGA playlist in the testing network. Incase if you need to have a look at it, please
make the SAC_LOKESH presentation as the default one and then run the Digital Signs Sidebar.

* I had uninstalled the Panopto from the NEXUS Mac Mini and had uninstalled all the extra components that got
installed on 23rd of Oct, 2008.
Worked hours:
Tue - 1.30PM - 7.15PM
Wed - 9.15AM - 2 PM
Thur - 12.10PM -3.20PM ,3.45PM-7PM




Rewa Tikekar:
10/21:
3:00PM to 6:00PM
7:00PM to 7:30PM
1. Installed new version for Recorder and Editor on laptop and did a quick check for recording with new recorder on
Laptop.
2. Continued with the Phase 1 Recordings with 2 machines. The upload for the previous lecture which was created
by lokesh is taking lot of time so we avoided uploading other lectures to minimize the error in time to upload due to
simultaneous uploads. We have just recorded the lectures and will upload once the first upload is done.
3. Capturing details in the logbook sent to you by Lokesh.

4. Comparing between previous video and video with new camera is concerned, DV mode was giving a good quality
picture compared to web only and HQC.

10/22:
Time: 9:00AM to 1:30PM
      4:00PM to 7:00PM
1.Attended the status update meeting.
2.Created 2 playlists to be played one after another in Nexus. I could see the switching between two playlists twice.
Though I could not figure out at what time stamp or what setting will it happen for sure.
During the test the schedule times given to playlists were 17:30 to 17:40 and second playlist was 17:40 to 17:50 as
we had discussed. I again changed the times keeping the same pattern for scheduling. At one point when I closed
and restarted the Nexus display window I could see the second playlist(I am not sure whether to coincide this with
closing and logging again into Nexus display).
10/23:
Time : 12:00 to 3:15PM
       3:40 to 7:45PM
1.Tried different sequences for scheduling.
  --> playing 2 playlists one after other.
  --> scheduling 2nd playlist just before the first one ends.
  I saw the playlists switching twice but it is not at all consistent.

2. Tried Panopto with 2 Canon cameras.
   --> on Mac mini.
   --> on dell with external.
   Recordings using fire wire for two cameras is not working with either of the above.
3. Created a playlist for SGA slides which are to be played from 27th to 31st
   These can be found under erbplasma presentation folder (to be set to default to view the playlist)
   The slides are saved under COMMON SHARED-->Announcements-->SGA.
   The playlist name is SGA Slide Playlist.
   The slide name is SGA1(this has a slide show of two given slides).


10/24:
Time: 2:00 to 3:30 PM
      5:00 to 6:30PM
1. Created a playlist for the slides Erbplasma.
2. Understood the playlists being played in production network.
3. Installed .NET 3.5 and NCS on Mac mini.
4. Performed initial tests on NCS and getting around with it.



 Jack M. Stein
Assistant Dean and Director of Engineering Computing The University of North Carolina at Charlotte
226 Cameron Research Building
9201 University City Boulevard
Charlotte, NC 28223
jmstein@uncc.edu
704-687-6625
412-848-6534 (cell)
704-687-2352 (fax)

				
DOCUMENT INFO
Shared By:
Categories:
Tags:
Stats:
views:0
posted:2/10/2013
language:Unknown
pages:16